Show that |βˆ«βˆ’2πœ‹2πœ‹ π‘₯ 2 sin8 (𝑒 π‘₯ ) 𝑑π‘₯| ≀ (16πœ‹ 3 )/3


1
Expert's answer
2021-09-27T04:10:59-0400

If f is integrable on [a,b], then |f| is integrable on [a,b] and


"|\int^b_a f(x)dx|\le \int^b_a |f|dx"


So:

"|\int^{2\pi}_{-2\pi} x^2sin^8(e^x)dx|\le \int^{2\pi}_{-2\pi} |x^2sin^8(e^x)|dx="

"=\int^{2\pi}_{-2\pi} |x^2||sin^8(e^x)|dx\le \int^{2\pi}_{-2\pi} |x^2|dx=\int^{2\pi}_{-2\pi} x^2dx=\frac{16\pi^3}{3}"